4kg of water at 50°C is poured on a large block of ice at 0°C. How much I've melts. Specific heat capacity of water= 4200J/Kg/°C. Latent heat of fusion of ice= 3.33x10^5J

1
Expert's answer
2022-04-01T08:41:17-0400

Given:

m1(H2O)=4 kg

T=50°C

T(ice)=0°C

Specific heat capacity of water; S= 4200J/Kg/°C

Latent heat of fusion of ice; Lf= 3,33x10^5J

m2(ice melted)=?

Solution:

Heat flow from higger temperature body to low temperature body

Heat lost by water=Heat gained by ice

Heat lost by water=m*S*ΔT=4*4200*50=840000 J

Heat gained by ice=m*Lf

m1*S*ΔT=m2*Lf

m2=m1*S*ΔT/Lf

m2=4*4200*50/3.33*105=2,52kg

Answer: 2,52 kg of ice melted